A binary liquid mixture is in equilibrium with its vapour at a temperature T=300K. The liquid mole fraction x1 of species 1 is 0.4 and the molar excess Gibbs energy is 200J/mol. the value of the universal gas constant is 8.314 J/mol/k and y1 denotes the liquid phase activity co-efficient of species.
1. If ln(y1)=0.09, find the value of ln(y2).
1
Expert's answer
2020-03-27T08:58:10-0400

Excess molar Gibbs (G) energy of a binary liquid mixture is expressed as:

G = RT*(x1*lny1 + x2*lny2)

It is given that x1 = 0.4, hence, x2 = 0.6.

Thus,

lny2 = (G/RT - x1*lny1)/x2 = (200/8.314/300 - 0.4*0.09)/0.6 = 0.0735